# Limits to Predicting Online Speech Using Large Language Models

### v2

@@ 4.4 User Tweets are Hard to Predict
In Figure 9 we present the distribution of the average model uncertainty on the tweets of our subjects. To gain an intuition on how hard it is to predict our subjects’ tweets relative to other accounts, we included three popular financial news accounts as comparison: YahooFinance, CNBC, and Stocktwits. Intuitively, predicting financial news and the stock market is hard Johnson et al. (2003), which should make them less predictable. However, Figure 9 reveals that most subjects in our pool are actually harder to predict than those news accounts. Of course, while the content of those tweets may be hard to predict, their style and vocabulary may not.
Finally, the absolute information content of users’ tweets is between 1.5-2 bits per character, depending on which model we use (see ’no context’ bars in Figure 2(c)). As a comparison, Shannon’s upper bound on the cross-entropy of the English language is 1.3 bits per character Shannon (1951), and recent estimates using neural language models say it is as low as 1.12 bits Takahashi and Tanaka-Ishii (2018). These results suggest that individual pieces of our online expression taken out of context are far from predictable using today’s LLMs. Provided with additional user context, only our largest model (Llama-2 70b) achieves comparable entropy, with an average of 1.1193 bits per character. While humans converge to 1.3 bits after only seeing $\sim$32 characters Moradi et al. (1998), language models typically need much longer context (see example from Appendix B.9).
@@ 5 Limitations
While negative log-likelihood is the most popular measure of model uncertainty, there are certain limitations to using this metric. One is that it is calculated exclusively on tokens contained in tweet $T$, and does not take into account any improvements on semantically equivalent text Kuhn et al. (2023). Another related issue is calibration Xiao et al. (2022). We also found the models to be highly sensitive to slight changes in prompting e.g. to how tweets were separated (App. A.3). Model uncertainty could also largely vary on the first token (App. B.6). Another important aspect is the length of input; NLL is commonly lower for longer pieces of text (App. B.3). While these points affect absolute numbers in our analysis, it does not affect our main statements about the relative comparisons between different models and contexts.
@@
Furthermore, we cannot rule out the possibility that some tweets in our corpus were part of the training dataset of the LLMs that we used. However, we believe the risk and severity of data contamination to be limited; we provide more detail as to why in Appendix A.2.
@@
Extending our analysis to different sub-communities and languages would be an interesting avenue for future work. Similarly, we welcome additional work into the external validity of our results. It is unclear to what extent our findings would translate to other social media platforms.
@@ 6 Discussion
We found that it is surprisingly difficult to predict online speech. The majority of our users’ tweets are less predictable than tweets about financial news. Even with additional context, prediction performs relatively poorly. Only our largest model with additional user context is able to approximate the estimated entropy rate of the English language. Further analysis reveals that the source of greatest uncertainty for base models are hashtags and @-mentions. In other words, what improves prediction the most with additional context are basic signals: specifically, guessing the correct hashtags and @-mentions. All in all, our findings suggest that despite the impressive capabilities of large language models in other areas, state-of-the-art models predict online speech rather poorly.
Another central question of our analysis was whether the predictive information inside peer tweets is enough to match (or even surpass) those of the user’s own tweets, as suggested by Bagrow et al. (2019). We believe that this is unlikely. Our results show that user context consistently outperforms peer context in a manner that is robust to model choice and evaluation method. Additional experiments suggest that whatever predictive information is inside the peer context, can also be found in the user context.
Our findings may also inform work on privacy and other ethical issues on social networks. For example, threats such as shadow profiling or impersonation on a global scale may not be as acute as some feared with state-of-the-art language models. However, these risks may still exist on a case-by-case basis (we demonstrate individual variability in Appendix B.2). Moreover, there may be other conceivable harms that don’t map cleanly to questions of predictability.
@@ A.1.1 Sampling stage
This collection phase resulted in a pool of 10M tweets, with 5M unique authors. We sampled our subjects from this pool of authors ($N=5102$ sample). Strictly speaking, our sample will be biased towards users that have 1) been more active during our initial collection period and 2) are more active users in general. This was the closest we could get to a random sample with the offered API endpoints. This practice follows Bagrow et al. (2019)’s method of randomly sampling users.
Twitter had roughly $\sim$500M active monthly users in 2023 Yaccarino (2023) – some of which may have been bot accounts Lee et al. (2011). To address this issue, we decided to filter out accounts that scored high (above 0.5 on a scale of 0 to 1) on the Bot-O-Meter bot detection API Sayyadiharikandeh et al. (2020). Bot-O-Meter uses classifies users on a scale from 0-1 based on 200 of their tweets. See Figure 10 for a distribution of these scores. We dropped users that had a score higher than 0.5. Additionally, we dropped users that had a high retweet ratio (more than $80\%$ of their tweets consisted of retweets). This is an additional measure to prevent bot accounts in our subject pool (bots are known for frequently retweeting content Yang et al. (2020); Gilani et al. (2017)) as well as a practical consideration since we only wanted to include non-retweets in our analysis.
@@ A.1.3 Preprocessing tweets
Before our experiments, we preprocessed our collected tweets by filtering out urls, deduplicated spaces and fixed some special character encodings. We found that urls were not relevant in analysing the predictability of (organic) online speech, while deduplication of spaces is a fairly common preprocessing step in NLP. These preprocessing choices were partially inspired by an open-source project called HuggingTweets Dayama (2022). For our experiments described in Section 4.2, we further filtered out @-mentions and hashtags.
@@
(b) Loss calculated on the evaluation set of 100 subjects ($25$k tweets). There is a lot more stochasticity, which in part is due to the smaller sample size and a less diverse pool of authors.
@@ A.2.1 Training data
GPT-2 was trained on WebText with content up to December 2017 Radford et al. (2019) ($95\%$ of our subjects’ tweets were authored after 2022), Falcon was trained on RefinedWeb which is is built using all CommonCrawl dumps until the 2023-06 one Penedo et al. (2023) (CommonCrawl typically does not contain snapshots of Twitter com (2024)) and the authors of the Llama models claim to have “made an effort to remove data from certain sites known to contain a high volume of personal information about private individuals.” Touvron et al. (2023); Dubey et al. (2024). Llama-2 has a knowledge cutoff at September 2022, while Llama-3 8B has seen data up to March 2023. Based on these facts we believe there is limited risk of data contamination.
@@ A.2.2 Tweet-tuning on 10M tweets
GPT-2-XL-tt
We finetuned all parameters for 1 epoch, using a constant learning rate of $5\mathrm{e}{-5}$, batch size 8 and fp16 mixed precision training on a single A-100 80GB GPU with an AdamW optimizer. We used the example finetuning script from the transformers library as our base222run_clm.py from here: https://github.com/huggingface/transformers/tree/main/examples/pytorch/language-modeling, where we kept most of the default training arguments.
@@ A.3 Prompting experiments
Let us denote a single tweet as $T=(t_{1},t_{2},...t_{m})$, where $t_{i}$ is a single token ($i=0...m$). A token is an item in the LLM’s vocabulary, and can be thought of as a collection of characters that frequently co-occur. Each tweet has a maximum of 280 characters [^4] This limit changed to 4000 characters on the 09. February 2023. and after tokenization most tweets have between 0-100 tokens (Fig. 20). We use a language model with parameters $\theta$ to predict token $t_{i}$ based on the preceding tokens $t_{<i}$. The model’s output will be a likelihood over all possible tokens in the model’s vocabulary, however we are only interested in the conditional probability of $t_{i}$: $p_{\theta}(t_{i}|t_{<i})$.
@@
This gives us a unit-free quantity to estimate the effect size. Following established practice Cohen (1988), we characterize effect sizes up to $0.2\sigma$ to be small, up to $0.8\sigma$ to be medium and and anything above that to be a large effect size.
@@ A.4 Finetuning experiments
We only ran this experiment on the tweet-tuned version of GPT-2 (GPT-2-XL-tt), because of the non-negligible amount of time and resources this experiment requires to run. Per subject there are 3 different types of context, plus 4 mixtures of contexts. That is, we finetuned GPT-2 a total of $7*5102=35714$ times. Together with tweet-tuning, it cost us approximately 1.6 e19 FLOPs to run this experiment.  [^6] As a comparison: it took OpenAI around e20-e21 FLOPs to finetune GPT-2 from scratch. While finetuning on larger models would certainly result in lower cross-entropy overall, we do not believe that it would change our main conclusions. This is supported by the high agreement on the ranking of users across models (Fig. 16) as well as the stability of our conclusions for our prompting experiments as we go up in model size.
@@ Appendix B Supplementary Results
Figure 17: Differences in predictability $\Delta^{c_{1}}_{c_{2}}$ using the Llama-2 model. We picked 10 random users and plot their predictability using different contexts (y axis). Comparing across contexts we get the difference in predictability: $\Delta(u_{i})^{c_{1}}_{c_{2}}$. While unpredictability goes down on average as we evaluate on more predictive contexts, there can be individual level variance that does not necessarily always follow this trend.
@@ B.5 Two control groups: social and temporal control
Bagrow et al. (2019) introduce two control groups in their experiment: a social and a temporal control. Social control includes tweets from 15 random users. Temporal control on the other hand, selects 250 tweets that were authored around the same time as the tweets from the peer context. See Figure 12 for an illustration of both. The rugplot on the top of Figure 13 shows the same, but on real data of a random subject. Figure 18 shows our main results including both controls, while Figure 22 shows the corresponding effect sizes.
@@ B.7 Most improved token due to random context: @-mention
We were interested in which tokens benefited most after including random (i.e. social) context. We selected tokens with >100 occurences, and ranked them based on how much their prediction improved on average. Table 1 shows that the ’@’ token benefits the most across all three models, indicating that @-mentions are locations of greatest improvement. See Table LABEL:tab:top_10_full for a full table on top-10 most improved token predictions (with all possible contexts).
@@ 10  Ton  Om  txt ihe
Table 2: Top ten tokens whose predictability went up the most (on average) after including tweets from some context (compared to no context). Tokens with $>100$ occurences were selected.

### v3

@@ 4.3 Results on Tweet-Tuned Models
Figure 8: Average loss curves with standard error for finetuning experiments on GPT-2-XL-$t$. For each subject ($N=5102$) we finetune on the specified context, and compute the cross-entropy loss on $\mathcal{T}_{u}^{\texttt{eval}}$. The plotted averages are computed over the loss curves of 1000 subjects.
@@ 4.4 User Tweets are Hard to Predict
In Figure 9 we present the distribution of the average model uncertainty on the tweets of our subjects. To gain an intuition on how hard it is to predict our subjects’ tweets relative to other accounts, we included three popular financial news accounts as comparison: YahooFinance, CNBC, and Stocktwits. Intuitively, predicting financial news and the stock market is hard johnson_financial_2003, which should make them less predictable. However, Figure 9 reveals that most subjects in our pool are actually harder to predict than those news accounts. Of course, while the content of those tweets may be hard to predict, their style and vocabulary may not.
@@
Finally, the absolute information content of users’ tweets is between 1.5-2 bits per character, depending on which model we use (see ’no context’ bars in Figure 2(c)). As a comparison, Shannon’s upper bound on the cross-entropy of the English language is 1.3 bits per character shannon_prediction_1951, and recent estimates using neural language models say it is as low as 1.12 bits takahashi_cross_2018. These results suggest that individual pieces of our online expression taken out of context are far from predictable using today’s LLMs. Provided with additional user context, only our largest model (Llama-2 70b) achieves comparable entropy, with an average of 1.1193 bits per character. While humans converge to 1.3 bits after only seeing $\sim$32 characters moradi_entropy_1998, language models typically need hundreds to thousands of tokens to converge (see Fig. 25 in Appendix B.9).
4.5 Results across groups
It is also important to analyze our results and see how they apply to different demographics. Are some groups more predictable than others? Do we discover the same relative relationships between contexts? For example, it may be possible that members of some group may be more likely to be profiled through their peers. While we did not have any demographic information that covered all our subjects, we came up with proxies that allowed us to analyze a subset of them. We analyzed our results by gender, location, and their intersection. The latter provided similar results as the first two and can be found in Appendix B.10.
We extracted 830 users’ preferred pronouns as a proxy for gender. This was done by searching for string matches in their profile descriptions. For example, profiles with “he/him” in their description were matched to ‘masculine’, “she/her” to ‘feminine’, and ones that contained a mix of both or other pronouns  [^3] The full list of pronouns we matched for is in Appendix B.10. were matched to the ‘diverse’ category. What we mean by a ‘proxy’ is that it merely implies the individual’s preferred linguistic gender, not necessarily their social gender devinney2022theories (i.e. she/her can be the preferred pronoun of both women and trans women). Results from Figure 10(a) show that LLMs achieve similar performance across all categories. Subjects across different groups seem to be equally predictable. Even the relative relationship between contexts is preserved — no group is significantly more likely to be profiled through their social circle.
We further identified 2221 users’ country based on the specified location in their profile. This was mostly done by using widely available geocoding services like Nominatim  [^4] https://nominatim.org/. We only analyzed the top-5 most common countries (United States, Great Britain, Canada, India and Nigeria). Compared to before, results in Figure 10(b) again show similar performance across countries except for Nigeria, where model uncertainty is significantly higher. This indicates that language models perform significantly worse on tweets belonging to Nigerian profiles. This may be in part due to Nigerian subjects speaking their own English dialects. Our results closely mirror prior work which have found that models perform significantly worse on certain dialects, such as African-American-English blodgett2016demographic.
(a) Results by gender. ‘masculine’: he/him (233), ‘feminine’: she/her (291), and ‘diverse’ (306): mix of feminine / masculine or other pronouns (like they/them).
(b) Results by location. Top-5 locations: ‘US’: United States (1555), ‘GB’: Great Britain (408), ‘CA’: Canada (124), ‘IN’: India (70) and ‘NG’: Nigeria (64).
Figure 10: Results across different demographic groups. Model: Llama-3-8B.
@@ 5 Limitations
While negative log-likelihood is the most popular measure of model uncertainty, there are certain limitations to using this metric. One is that it is calculated exclusively on tokens contained in tweet $T$, and does not take into account any improvements on semantically equivalent text kuhn2023semantic. Another related issue is calibration xiao2022uncertainty. We also found the models to be highly sensitive to slight changes in prompting. See the Appendix how the separator token between tweets A.3, the first token B.6), and tweet length B.3 affect model uncertainty. While these points might affect absolute numbers in our analysis, it does not affect our main statements about the relative comparisons between different models and contexts.
Despite these limitations, there are compelling reasons to use negative log-likelihood. It is the metric typically optimized during language model training (via cross-entropy loss) and is widely used in NLP as a proxy for downstream model performance. The core assumption is that differences in loss between models reflect differences in their underlying capabilities saunshi2020mathematical. Directly evaluating certain capabilities – such as the ability of large language models to impersonate individuals – can raise ethical concerns, making NLL a valuable alternative.
@@
Furthermore, we cannot rule out the possibility that some tweets in our corpus were part of the training dataset of the LLMs that we used. Contamination may result in lower model uncertainty. However, we believe the risk and severity of data contamination to be limited; we provide more detail as to why in Appendix A.2.
@@
Our analysis is constrained to the English-speaking population. How our results extend to other languages and dialects would be an extremely valuable avenue for future work. Similarly, a deeper analysis into how minorities and individuals are affected would be necessary. While Twitter is not the only widely used social media platform, it is one of the primary text-focused ones. Our findings may extend to other similar platforms (such as Mastodon, BlueSky and Threads), but they may apply less to platforms that focus more on sharing image/video content (such as Instagram and TikTok).
@@ 6 Discussion
Our main finding is that online speech is surprisingly hard to predict. Most users’ tweets are less predictable than tweets about financial news, and even with additional context, predictability remains relatively low. Only our largest model with additional user context is able to approximate the estimated entropy rate of the English language. We also found that additional context improves the prediction of basic signals: specifically, guessing the correct hashtags and @-mentions. All in all, our findings suggest that despite the impressive capabilities of large language models in other areas, state-of-the-art models predict speech rather poorly.
Similarly, our results indicate that it matters greatly what context we use for prediction. We questioned whether the predictive information inside peer tweets is enough to match (or even surpass) those of the user’s own tweets, as suggested by bagrow_information_2019. We believe that this is unlikely. Our results show that user context consistently outperforms peer context in a manner that is robust to model choice and evaluation method. Additional experiments suggest that whatever predictive information is inside the peer context, can also be found in the user context.
To summarize, we have shown that predicting online speech is rather difficult and that it depends a lot on what information is used as context. Our findings are robust across methods (prompting vs. finetuning) and show consistent trends across different demographic groups (gender and location).
Studying the predictability of online speech may inform work on privacy and other ethical issues on social networks. For example, threats such as shadow profiling or impersonation on a global scale may not be as acute as some feared with state-of-the-art language models. However, we do find variability on the level of the individual users (see Appendix B.2). Moreover, there may be other conceivable harms that don’t map cleanly to questions of predictability. Future work could attempt to build a bridge between model uncertainty and downstream capabilities that affect the individual, ensuring continued scrutiny as language models inevitably evolve.
@@ A.1.1 Sampling stage
This collection phase resulted in a pool of 10M tweets, with 5M unique authors. We sampled our subjects from this pool of authors ($N=5102$ sample). Strictly speaking, our sample will be biased towards users that have 1) been more active during our initial collection period and 2) are more active users in general. This was the closest we could get to a random sample with the offered API endpoints. This practice follows bagrow_information_2019’s method of randomly sampling users.
Twitter had roughly $\sim$500M active monthly users in 2023 linda_twitter_2023 – some of which may have been bot accounts lee_seven_2011. To address this issue, we decided to filter out accounts that scored high (above 0.5 on a scale of 0 to 1) on the Bot-O-Meter bot detection API sayyadiharikandeh_detection_2020. Bot-O-Meter uses classifies users on a scale from 0-1 based on 200 of their tweets. See Figure 11 for a distribution of these scores. We dropped users that had a score higher than 0.5. Additionally, we dropped users that had a high retweet ratio (more than $80\%$ of their tweets consisted of retweets). This is an additional measure to prevent bot accounts in our subject pool (bots are known for frequently retweeting content yang_scalable_2020; gilani_bots_2017) as well as a practical consideration since we only wanted to include non-retweets in our analysis.
@@ A.1.3 Preprocessing tweets
Before our experiments, we preprocessed our collected tweets by filtering out urls, deduplicated spaces and fixed some special character encodings. We found that urls were not relevant in analysing the predictability of (organic) online speech, while deduplication of spaces is a fairly common preprocessing step in NLP. These preprocessing choices were partially inspired by an open-source project called HuggingTweets dayama_huggingtweets_2022. For our experiments described in Section 4.2, we further filtered out @-mentions and hashtags.
@@
(b) Loss calculated on the evaluation set of 100 subjects (2525k tweets). There is a lot more stochasticity, which in part is due to the smaller sample size and a less diverse pool of authors.
@@ A.2.1 Training data
GPT-2 was trained on WebText with content up to December 2017 radford2019language ($95\%$ of our subjects’ tweets were authored after 2022), Falcon was trained on RefinedWeb which is is built using all CommonCrawl dumps until the 2023-06 one penedo_refinedweb_2023 (CommonCrawl typically does not contain snapshots of Twitter commoncrawl_statistics_2024) and the authors of the Llama models claim to have “made an effort to remove data from certain sites known to contain a high volume of personal information about private individuals.” touvron_llama_2023; dubey2024llama. Llama-2 has a knowledge cutoff at September 2022, while Llama-3 8B has seen data up to March 2023. Based on these facts we believe there is limited risk of data contamination.
@@ A.2.2 Tweet-tuning on 10M tweets
GPT-2-XL-$t$
We finetuned all parameters for 1 epoch, using a constant learning rate of $5\mathrm{e}{-5}$, batch size 8 and fp16 mixed precision training on a single A-100 80GB GPU with an AdamW optimizer. We used the example finetuning script from the transformers library as our base555run_clm.py from here: https://github.com/huggingface/transformers/tree/main/examples/pytorch/language-modeling, where we kept most of the default training arguments.
@@ A.3 Prompting experiments
Let us denote a single tweet as $T=(t_{1},t_{2},...t_{m})$, where $t_{i}$ is a single token (i=0mi=0...m). A token is an item in the LLM’s vocabulary, and can be thought of as a collection of characters that frequently co-occur. Each tweet has a maximum of 280 characters [^7] This limit changed to 4000 characters on the 09. February 2023. and after tokenization most tweets have between 0-100 tokens (Fig. 21). We use a language model with parameters $\theta$ to predict token $t_{i}$ based on the preceding tokens $t_{<i}$. The model’s output will be a likelihood over all possible tokens in the model’s vocabulary, however we are only interested in the conditional probability of $t_{i}$: $p_{\theta}(t_{i}|t_{<i})$.
@@
This gives us a unit-free quantity to estimate the effect size. Following established practice cohen_statistical_1988, we characterize effect sizes up to $0.2\sigma$ to be small, up to $0.8\sigma$ to be medium and and anything above that to be a large effect size.
@@ A.4 Finetuning experiments
We only ran this experiment on the tweet-tuned version of GPT-2 (GPT-2-XL-tt), because of the non-negligible amount of time and resources this experiment requires to run. Per subject there are 3 different types of context, plus 4 mixtures of contexts. That is, we finetuned GPT-2 a total of 75102=357147*5102=35714 times. Together with tweet-tuning, it cost us approximately 1.6 e19 FLOPs to run this experiment.  [^9] As a comparison: it took OpenAI around e20-e21 FLOPs to finetune GPT-2 from scratch. While finetuning on larger models would certainly result in lower cross-entropy overall, we do not believe that it would change our main conclusions. This is supported by the high agreement on the ranking of users across models (Fig. 17) as well as the stability of our conclusions for our prompting experiments as we go up in model size.
@@ Appendix B Supplementary Results
Figure 18: Differences in predictability $\Delta^{c_{1}}_{c_{2}}$ using the Llama-2 model. We picked 10 random users and plot their predictability using different contexts (y axis). Comparing across contexts we get the difference in predictability: $\Delta(u_{i})^{c_{1}}_{c_{2}}$. While model uncertainty goes down on average as we evaluate on more predictive contexts, there can be individual level variance that does not necessarily always follow this trend.
@@ B.5 Two control groups: social and temporal control
bagrow_information_2019 introduce two control groups in their experiment: a social and a temporal control. Social control includes tweets from 15 random users. Temporal control on the other hand, selects 250 tweets that were authored around the same time as the tweets from the peer context. See Figure 13 for an illustration of both. The rugplot on the top of Figure 14 shows the same, but on real data of a random subject. Figure 19 shows our main results including both controls, while Figure 23 shows the corresponding effect sizes.
@@ B.7 Most improved token due to random context: @-mention
We were interested in which tokens benefited most after including random (i.e. social) context. We selected tokens with >100 occurences, and ranked them based on how much their prediction improved on average. Table 1 shows that the ’@’ token benefits the most across all three models, indicating that @-mentions are locations of greatest improvement. See Table 2 for a full table on top-10 most improved token predictions (with all possible contexts).
@@ B.9 Negative log-likelihood convergence
B.10 Results across groups
B.10.1 Gender
We applied the following regular expressions on users’ names, descriptions and location to match their preferred pronouns. We used the ‘feminine’, ‘masculine’ and ‘diverse’ categories. Case was ignored.
The regex for the ‘feminine’ category:
[…]
The regex for the ‘masculine’ category:
[…]
The regex for the ‘diverse’ category:
[…]
Subjects that matched for pronouns in both the ‘feminine’ and ‘masculine’ category were also put in the ‘diverse’ group.
B.10.2 Location
Users’ location was determined using geolocating services like Nominatim. We used python libraries like geopy and geotext to extract the country code of the location. Another method was to parse flag emojis and convert them to the corresponding two-letter country code. The input was the specified location in users’ profiles.
B.10.3 Intersection of gender and US
We also had enough data to analyze the intersection of US subjects and the gender category. Results are in Fig. 26. We see similar predictability and relative relationships in these subcategories as well.
Figure 26: Intersection of gender and US. Categories: US & masculine (78), US & feminine (69) and US & diverse (81). Model: Llama-3-8B
@@ 10  Ton  Om  txt ihe
Table 2: Top ten tokens whose predictability went up the most (on average) after including tweets from some context (compared to no context). Tokens with >100>100 occurences were selected.